2+ 5 + 1 + 2 +2) Consider an insurance market with insurance firms being competitive...

  1. 2+ 5 + 1 + 2 +2) Consider an insurance market with insurance firms being competitive and risk neutral (Zero expected profits in equilibrium) and a risk averse customer with a von-Neumann Morgenstern utility function . The customer is born with a wealth of $150. There is a probability of an accident p=0.3 in which case the customer will suffer a damage of $50.

An insurance contract is , where premium paid in the no accident state, and = net payout received in case of accident

  1. What is the customer’s utility without insurance?
  2. If all Zero profit insurance contracts are offered, write down the customer’s utility maximization problem. What contract would she choose? What will be the premium? What is the payout?
  3. Is the customer happier with insurance? Show why or why not.
  4. If there was another customer with p =0.5, (a high-risk customer) what contract will she choose? (assume everything else for this new customer is the same as the old one, and the customers are identifiable). Show your work.
  5. Suppose both contracts (for part (ii) and (iv)) are being offered by insurance companies, and all customers are free to choose whatever contract they prefer. Which one will the high type customer choose? Show your work.
Answer #1

Solution:

o 0 G+ Good state XG = 150, Pg 2 (IP) 2:4, X8 = 150-50 =100, PR =P=03 B, ad loss stalē bot EU (NO Insurance) = .74 (150) to 3

113 DATE: PAGE: P FOC P Now EU (with Insurance) EU = 04 (XG-1) 13 + 03 (XB +_z) 113 04 (150- ov. 7 o 3 (100 + (1-P) (1) aeu ¥

1596354189133_IMG_20200802_131023.png

NOTE : I HOPE YOUR HAPPY WITH MY ANSWER PLEASE GIVE ME "LIKE"